### v1

# Unified quantification of entanglement and magic in information scrambling and their trade-off relationship

Abstract
Information scrambling is a phenomenon observed in a wide range of quantum systems, in which initially localized information becomes distributed across the entire system. Clarifying the essence of this phenomenon is made possible by quantum resource theory, which provides an axiomatic framework for quantifying various forms of quantumness. Within this framework, information scrambling can be classified into two distinct types: entanglement scrambling and magic scrambling. Although each has been analyzed individually through their respective resource theories, such separate treatments fail to capture the inherent trade-off between them. In this study, we introduce a unified measure of the two types of scrambling, grounded in resource-theoretic principles. This unified approach reveals a rigorous trade-off relationship between entanglement and magic scrambling, as the exact maximum value of the proposed measure can be derived analytically. Furthermore, we quantify the scrambling capability of unitary transformations in terms of their ability to amplify this measure.
I Introduction
Quantum systems exhibit a range of non-classical and counterintuitive behaviors. Understanding the origins of such quantumness remains one of the most fundamental challenges in quantum mechanics. Quantum resource theories provide a powerful framework to capture and formalize the essential features of quantum systems (see Ref. [1] for a review). In a typical resource-theoretic framework, a fixed set of operations—known as free operations—is specified in advance, and any property that cannot be generated using only these operations is regarded as a resource. To quantify the resource, a resource measure (a function that assigns a numerical value to the resource) is introduced, typically satisfying several core properties such as faithfulness and monotonicity. This formalism was originally developed to quantify entanglement. Since then, a variety of quantum features, including coherence [2, 3, 4, 5], superposition [6], magic [7], uncomplexity [8], and thermodynamic nonequilibrium [9, 10], have been studied within this framework. Constructing a resource theory for a given quantum property enables both its characterization and quantitative analysis.
Chaos has been extensively studied in mathematics and physics. One of the defining features of chaos is the butterfly effect, in which a slight change in initial conditions can lead to drastically different outcomes. In physical systems, chaotic behavior is observed in many-body systems with complex interactions, as well as in black holes. In quantum systems, chaos gives rise to quantum information scrambling, a process in which information initially localized in a small subsystem becomes distributed throughout the entire system. This scrambling has recently been linked to a wide range of fields, including quantum error correction [11], quantum machine learning [12, 13, 14, 15], or Clifford decoder [16]. Viewing information scrambling as a resource provides a powerful framework for understanding and quantifying this phenomenon.
Information scrambling has been quantified using various measures, such as the out-of-time-ordered correlator (OTOC) [17, 18, 19], the entanglement entropy [20, 21], and the stabilizer Rényi entropy [22, 23, 24]. Since these measures capture different aspects of scrambling, ambiguity remains as to what constitutes a highly scrambling transformation. Recently, a rigorous framework for the resource theory of information scrambling has been proposed [25, 26]. In the resource-theoretic context, scrambling is categorized into two distinct types based on underlying mechanisms: entanglement scrambling and magic scrambling (see FIG. 1). While these names are not yet standardized, they suggest that each type of scrambling can be understood as a quantum resource. Entanglement scrambling refers to transformations that increase the number of non-identity components in an operator’s Pauli basis expansion. It can be quantified by measures such as Pauli growth [25] and circuit sensitivity [26], which behave as valid resource monotones when non-entangling unitaries (i.e., single-qubit unitaries and swap operations) are taken as free operations. In contrast, magic scrambling corresponds to transformations that increase the number of terms required in the Pauli basis expansion. It is quantified by measures such as OTOC magic [25] and magic power [26], which are valid resource monotones under Clifford unitaries as the set of free operation. These resource measures satisfy several fundamental properties that justify their interpretation as resource monotones. Grounded in this rigorous framework, they characterize each type of scrambling individually.
In this study, we reconstruct the resource-theoretic framework for information scrambling. We define measures that treat both entanglement scrambling and magic scrambling in parallel. Furthermore, we focus on the fact that entanglement scrambling and magic scrambling cannot be simultaneously maximized. To clarify the trade-off relationship between them, we introduce a new quantity called operator complexity, which captures both types of scrambling in a unified way. As our most significant result, we analytically derive the exact maximum value of the operator complexity—revealing a fundamental limitation inherent to information scrambling. In addition, as a measure of the ability to amplify the operator complexity, we introduce a unified scrambling measure. These quantities satisfy several essential properties as resource monotones, thereby ensuring the soundness of our framework. The proofs for most propositions are provided in the appendix.
Figure 1: Schematic diagrams illustrating two types of information scrambling. (a) Entanglement scrambling corresponds to transformations that increase the number of non-identity operators in Pauli strings. (b) Magic scrambling corresponds to transformations that increase the number of terms required in the Pauli basis expansion.
II Scrambling of operators
Information scrambling can be defined as the ability of unitary transformations to amplify the complexity of operators. Entanglement scrambling and magic scrambling, which are different types of scrambling arising from distinct mechanisms, can be quantified using different measures of the complexity of operators. In this section, we introduce two measures of operators that correspond to each scrambling in the context of resource theory. Furthermore, we define a unified measure that takes into account both scrambling and derive its rigorous maximum value. This clarifies a trade-off between the two types of scrambling.
We consider an $n$-qubit system in the $2^{n}$-dimensional Hilbert space $\mathcal{H}$. Let $O$ be a normalized operator acting on $\mathcal{H}$, which can be expanded in the Pauli basis as $O=\sum_{i}c_{i}P_{i}$. Here, $\{P_{i}\}_{i}$ denotes the subset of the Pauli group on $\mathcal{H}$ consisting of elements with coefficients equal to +1+1. The coefficients satisfy the condition $\sum_{i}|c_{i}|^{2}=1$, ensuring that the norm of $O$ is normalized to unity.
Firstly, we focus on entanglement scrambling. As illustrated in FIG. 1, entanglement scrambling corresponds to transformations that increase the number of non-identity operators in the Pauli strings. To quantify this spreading, we introduce the average Pauli weight [25] of the normalized operator $O$, defined as
$W(O)\coloneqq\sum_{i}|c_{i}|^{2}W(P_{i}).$ (1)
For any Pauli operator $P$, $W(P)$ denotes the number of non-identity operators composing the Pauli string. For any normalized operator $O$, the average Pauli weight has the following basic properties:
  1. 1.
    Faithfulness: $W(O)\geq 0$, and $W(O)=0$ if and only if $O$ is the identity operator.
  2. $\tilde{1}$.
    Faithfulness (traceless case): $W(O)\geq 1$ for any traceless $O$, and $W(O)=1$ if and only if $O$ can be expanded using only weight-1 Pauli operators.
  3. 2.
    Invariance: $W(U^{\dagger}OU)=W(O)$ for any non-entangling unitary $U$.
  4. 3.
    Additivity: $W(O_{1}\otimes O_{2})=W(O_{1})+W(O_{2})$ for any normalized operators $O_{1}$ and $O_{2}$.
Here, a non-entangling unitary refers to a unitary that can be generated from single-qubit unitaries and swap operations.
Next, we consider the quantification of magic scrambling, which corresponds to an increase in the number of terms in the Pauli basis expansion of the operator (see FIG. 1). We introduce a measure of the complexity of the Pauli basis expansion, referred to as the quantum Fourier entropy [26], defined as
$S_{a}(O)\coloneqq-\sum_{i}|c_{i}|^{2}\log_{a}|c_{i}|^{2}.$ (2)
Here, we take the base of the logarithm as a parameter $a$. If $a$ is omitted, we set $a=4$ so that the maximum value of the quantum Fourier entropy becomes $n$. This choice aligns the ranges of $W(O)$ and $S(O)$, enabling us to treat the two types of scrambling on equal footing. For any normalized operator $O$, the quantum Fourier entropy satisfies the following properties:
  1. 1.
    Faithfulness: $S_{a}(O)\geq 0$, and $S_{a}(O)=0$ if and only if $O$ is a Pauli operator.
  2. 2.
    Invariance: $S_{a}(U^{\dagger}OU)=S_{a}(O)$ for any Clifford unitary $U$.
  3. 3.
    Additivity: $S_{a}(O_{1}\otimes O_{2})=S_{a}(O_{1})+S_{a}(O_{2})$ for any normalized operators $O_{1}$ and $O_{2}$.
$W(O)$ and $S_{a}(O)$ serve as resource measures under the respective resource theories where non-entangling unitaries and Clifford unitaries are considered free operations. Here, we define a resource-theoretic measure of the complexity of operators. Since this function captures both types of scrambling, it enables a unified treatment of information scrambling:
$R_{a}(O)\coloneqq W(O)+S_{a}(O)=\sum_{i}|c_{i}|^{2}\log_{a}\frac{a^{W(P_{i})}}{|c_{i}|^{2}}.$ (3)
We refer to this measure as operator complexity. The main contribution of this study is the introduction and analysis of the operator complexity. $R_{a}(O)$ quantifies the extent to which a normalized operator simultaneously exhibits entanglement and magic scrambling. Since $W(O)$ and $S(O)$ share the same range of values, when the parameter $a$ is set to 4, Eq. (3) treats both types of scrambling equally. For any normalized operator $O$, the operator complexity satisfies the following properties:
  1. 1.
    Faithfulness: $R_{a}(O)\geq 0$, and $R_{a}(O)=0$ if and only if $O$ is the identity operator.
  2. $\tilde{1}$.
    Faithfulness (traceless case): $R_{a}(O)\geq 1$ for any traceless $O$, and $R_{a}(O)=1$ if and only if $O$ is a weight-1 Pauli operator.
  3. 2.
    Invariance: $R_{a}(U^{\dagger}OU)=R_{a}(O)$ for any non-scrambling unitary $U$.
  4. 3.
    Additivity: $R_{a}(O_{1}\otimes O_{2})=R_{a}(O_{1})+R_{a}(O_{2})$ for any normalized operators $O_{1}$ and $O_{2}$.
Here, non-scrambling unitaries are defined as the intersection of non-entangling unitaries and Clifford unitaries; that is, any such unitary can be generated from single-qubit Clifford unitaries and swap operations. These properties ensure that $R_{a}(O)$ is a valid resource monotone. They can be readily derived from the properties of the average Pauli weight and the quantum Fourier entropy.
$W(O)$ and $S_{a}(O)$ cannot be simultaneously maximized; that is, there exists a trade-off between entanglement scrambling and magic scrambling. Here, we derive the exact maximum value of the operator complexity. This result reveals a fundamental limitation inherent in information scrambling.
Since the value of the operator complexity is independent of the complex phases of the coefficients, we represent the normalized operator $O$ as $O=\sum_{i}\sqrt{p_{i}}P_{i}$. Under the constraints $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$, we solve the following optimization problem:
$\mathrm{maximize}\;\;R_{a}(O)=\sum_{i}p_{i}\left\{W(P_{i})-\log_{a}p_{i}\right\}.$ (4)
We provide the solution to this problem in the appendix. As a result, the closed-form expression of the operator that maximizes $R_{a}(O)$ is derived as
$O_{\mathrm{Max}}=\sum_{w=0}^{n}\sum_{W(P_{i})=w}e^{i\theta_{i}}\sqrt{\frac{a^{w}}{(1+3a)^{n}}}P_{i},$ (5)
where $\{\theta_{i}\}_{i}$ are arbitrary phases. This operator exhibits optimal spreading from both scrambling perspectives. Furthermore, the values of each measure obtained from $O_{\mathrm{Max}}$ can be computed as follows:
$W(O_{\mathrm{Max}})=\frac{3a}{1+3a}n,$ (6)
$S_{a}(O_{\mathrm{Max}})=n\log_{a}(1+3a)-\frac{3a}{1+3a}n,$ (7)
$R_{a}(O_{\mathrm{Max}})=n\log_{a}(1+3a).$ (8)
See the appendix for the details of the calculation. In particular, when the parameter $a$ is set to 4, we obtain
$R(O)\leq R(O_{\mathrm{Max}})\simeq 1.85n<2n.$ (9)
This is a key finding that stems from the operator complexity. Since the maximum values of $W(O)$ and $S(O)$ are both $n$, the operator complexity is clearly bounded above by $2n$. However, Eq. (9) reveals the upper bound strictly smaller than $2n$. This bound represents the exact maximum of the operator complexity, reflecting a fundamental constraint inherent in scrambling dynamics. Similarly, for general values of the parameter $a$, we obtain the following inequality:
$R_{a}(O)\leq R_{a}(O_{\mathrm{Max}})=n\log_{a}(1+3a)<(1+\lambda)n,$ (10)
where $\lambda=\log_{a}4$. The quantity $(1+\lambda)n$ is a trivial upper bound, since $R_{a}(O)=W(O)+\lambda S(O)$. Eq. (10) also captures the fundamental limitation inherent in the two types of scrambling.
In FIG. 2, we map the several representative operators on the $W$-$S$ plane. First, any weight-$n$ Pauli operator—for example, $O_{1}=Z_{1}\otimes\cdots\otimes Z_{n}$—gives the maximum value of $W(O)$, while $S_{a}(O)$ remains zero. Linear combinations of weight-$n$ Pauli operators also attain the maximum value of $W(O)$, and although they yield nonzero values of $S_{a}(O)$, they do not maximize it. For instance, the operator $O_{2}$ shown in FIG. 2 maximizes $W(O)$ and achieves $S_{a}(O_{2})=n\log_{a}3$, which is smaller than the maximum value of $S_{a}(O)$. In contrast, the uniform linear combination of all Pauli basis elements with equal amplitude, $O_{3}=1/2^{n}\sum_{i}P_{i}$, maximizes $S_{a}(O)$, while its average Pauli weight is not maximized: $W(O_{3})=3n/4$. Most importantly, $O_{\mathrm{Max}}$ is optimal with respect to the operator complexity $R_{a}(O)$, even though it does not individually maximize $W(O)$ or $S_{a}(O)$. The exact location of $O_{\mathrm{Max}}$ as a function of the parameter $a$ is also illustrated in FIG 2. In the limit $a\rightarrow 1$, $O_{\mathrm{Max}}$ converges to $O_{3}$, whereas in the limit $a\rightarrow\infty$, it converges to $O_{2}$.
Figure 2: Overview of the $W$-$S$ plane for arbitrary linear operators. A key observation is the trade-off relation between the two types of scrambling, illustrated by the solid red curve. When $W(O)$ reaches its maximum, $S_{a}(O)$ cannot, and vice versa—when $S_{a}(O)$ is maximized, $W(O)$ takes a specific, non-maximal value. The dashed curve provides a rough sketch of the general trend (possibly imprecise).
III Scrambling of unitaries
The measures introduced in the previous section allow us to quantify information scrambling exhibited by unitary transformations. The average Pauli weight and the quantum Fourier entropy are associated with entanglement scrambling and magic scrambling, respectively. Furthermore, the operator complexity provides a unified framework for quantifying information scrambling.
We define the following measures corresponding to each type of scrambling, which we refer to as entanglement growth, magic growth, and complexity growth, respectively:
$G_{E}(U)\coloneqq\max_{\|O\|_{2}=1}\left\{W(U^{\dagger}OU)-W(O)\right\},$ (11)
$G_{M}(U)\coloneqq\max_{\|O\|_{2}=1}\left\{S(U^{\dagger}OU)-S(O)\right\},$ (12)
$G(U)\coloneqq\max_{\|O\|_{2}=1}\left\{R(U^{\dagger}OU)-R(O)\right\}.$ (13)
The definitions presented here are inspired by the idea discussed in Refs. [25, 26]. Here, we set the parameter $a$ to 4, but the definition extends naturally to general values. These functions satisfy the following fundamental properties:
  1. 1.
    Faithfulness: $G_{(E,M)}(U)\geq 0$, and $G_{(E,M)}(U)=0$ if and only if $U$ is a free unitary.
  2. 2.
    Invariance: $G_{(E,M)}(V_{1}UV_{2})=G_{(E,M)}(U)$ for any free unitaries $V_{1},V_{2}$.
  3. 3.
    Subadditivity: $G_{(E,M)}(U_{1}\otimes U_{2})\leq G_{(E,M)}(U_{1})+G_{(E,M)}(U_{2})$ for any unitaries $U_{1}$ and $U_{2}$.
Non-entangling unitaries, Clifford unitaries, and non-scrambling unitaries constitute the set of free unitaries for $G_{E}(U)$, $G_{M}(U)$, and $G(U)$, respectively. Thus, these serve as valid measures from the perspective of resource theory.
Here, we provide an overview of the resource-theoretic framework for information scrambling (see TABLE 1.). We introduced three distinct operator-based measures; $W(O)$, $S(O)$, and $R(O)$. In particular, $R(O)$ captures both entanglement and magic scrambling. Based on the ability of unitary transformations to amplify these scrambling-related features, we defined corresponding measures of information scrambling. These measures satisfy several fundamental properties, allowing them to be treated as resources. For each type of scrambling, an appropriate set of free unitaries is identified. Taken together, the framework exhibits a parallel structure across the different scrambling types.
Finally, we introduce an additional measure of complexity growth:
$\tilde{G}(U)\coloneqq\max_{W(O)=1,S(O)=0}\left\{R(U^{\dagger}OU)-R(O)\right\}.$ (14)
This measure satisfies the following properties:
  1. 1.
    Faithfulness: $\tilde{G}(U)\geq 0$, and $\tilde{G}(U)=0$ if and only if $U$ is a non-scrambling unitary.
  2. 2.
    Invariance: $\tilde{G}(V_{1}UV_{2})=\tilde{G}(U)$ for any non-scrambling unitaries $V_{1},V_{2}$.
  3. 3.
    Maxitivity: $\tilde{G}(U_{1}\otimes U_{2})=\max\left\{\tilde{G}(U_{1}),\tilde{G}(U_{2})\right\}$ for any unitaries $U_{1}$ and $U_{2}$.
These support the validity of $\tilde{G}(U)$ as a resource monotone. In particular, maxitivity—a stronger property than subadditivity—is a distinctive feature of $\tilde{G}(U)$. $\tilde{G}(U)$ can be regarded as a more suitable measure for quantifying the information scrambling of a localized unitary. For instance, consider a localized unitary of the form $U=U_{1}\otimes I_{2}$. In this case, $\tilde{G}(U)=\tilde{G}(U_{1})$ can accurately evaluate the scrambling capability, whereas $G(U)\leq G(U_{1})$ may dilute it. In addition, since the set of operators satisfying $W(O)=1$ and $S(O)=0$ coincides with the set of weight-1 Pauli operators, $\tilde{G}(U)$ has the advantage of being numerically computable. Taken together, these observations suggest that $\tilde{G}(U)$ is a promising candidate for a unified measure of information scrambling.
Table 1: Overview of the measures for each type of information scrambling. A parallel structure emerges based on the resource-theoretic framework, with each type of scrambling associated with its own set of free operations.
  Operator measure   Unitary measure   Free unitary
  $W(O)=\sum_{i}|c_{i}|^{2}W(P_{i})$   $G_{E}(U)\coloneqq\max\left[W(U^{\dagger}OU)-W(O)\right]$   Non-entangling unitary
  $S(O)=-\sum_{i}|c_{i}|^{2}\log|c_{i}|^{2}$   $G_{M}(U)\coloneqq\max\left[S(U^{\dagger}OU)-S(O)\right]$   Clifford unitary
  $R(O)=\sum_{i}|c_{i}|^{2}(W(P_{i})-\log|c_{i}|^{2})$   $G(U)\coloneqq\max\left[R(U^{\dagger}OU)-R(O)\right]$   Non-scrambling unitary
IV Conclusion
In this study, we constructed a unified framework of resource theory for information scrambling. In particular, we introduced the concept of operator complexity, which accounts both entanglement scrambling and magic scrambling. As a main result, we derived the exact maximum value of this measure, revealing a trade-off relationship inherent in information scrambling. In addition, we quantified the scrambling capability intrinsic to unitary transformations as the amplification of such measures. For the complexity growth, we also introduced an alternative definition, which serves as another plausible resource measure. These measures exhibit a parallel structure and satisfy the properties required of resource monotones.
We anticipate that this work will contribute to a deeper understanding of the essence of information scrambling. Moreover, since our framework is built on measures for general linear operators, it is expected to extend naturally to more general settings, such as non-unitary channels. We hope that this study will inspire further developments across a broad range topics in quantum physics.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP24KJ0960, JP23K24915 and JP24K03008.
This appendix provides detailed proofs of the propositions presented in the main text.
Appendix A Average Pauli weight
We present the basic properties of the average Pauli weight $W(O)$ and provide reconstructed proofs based on Ref. [25].
Proof: Faithfulness.
By definition of the average Pauli weight, $W(O)\geq 0$ holds trivially. Equality holds if and only if $O$ is the identity operator. If $O$ is traceless, it contains no identity term in its Pauli expansion. Thus, for any traceless operator, $W(O)\geq 1$ holds, and $W(O)=1$ if and only if the Pauli expansion of $O$ consists solely of weight-1 Pauli operators. ∎
Proof: Invariance..
We prove the invariance of the average Pauli weight: for nay operator $O$, if $U$ is a non-entangling unitary, then $W(U^{\dagger}OU)=W(O)$. Here, a non-entangling unitary refers to any unitary that can be generated by single-qubit unitaries and swap operations.
In this proof, we expand the operator as $O=\sum_{\vec{a}}c_{\vec{a}}P_{\vec{a}}$, where the index $j$ denotes the position of the local Pauli operator. The average Pauli weight can then be computed as follows:
$W(O)$ $=\sum_{j}\sum_{\vec{a}(P_{j}\neq I)}|c_{\vec{a}}|^{2}$
$=\sum_{j}\frac{1}{d^{2n}}\sum_{\vec{a}(P_{j}\neq I)}\left|\mathrm{Tr}[OP_{\vec{a}}]\right|^{2}$
$=\sum_{j}\left[\frac{1}{d^{2n}}\sum_{\vec{a}}\left|\mathrm{Tr}[OP_{\vec{a}}]\right|^{2}-\frac{1}{d^{2n}}\sum_{\vec{a}(P_{j}=I)}\left|\mathrm{Tr}[OP_{\vec{a}}]\right|^{2}\right]$
$=\sum_{j}\left[\frac{1}{d^{n}}\mathrm{Tr}[O^{\dagger}O]-\frac{1}{d^{n+1}}\mathrm{Tr}\left[(\mathrm{Tr}_{j}[O])^{\dagger}(\mathrm{Tr}_{j}[O])\right]\right].$ (15)
Here, $d=2$ and $\mathrm{Tr}_{j}$ denotes the partial trace over the $j$th qubit. We note that the partial trace operation removes all terms except those corresponding to Pauli strings with $P_{j}=I$. When $U$ is generated as a product of single-qubit unitaries, the same calculation yields the following:
$W(U^{\dagger}OU)$ $=\sum_{j}\left[\frac{1}{d^{n}}\mathrm{Tr}[(U^{\dagger}OU)^{\dagger}(U^{\dagger}OU)]-\frac{1}{d^{n+1}}\mathrm{Tr}\left[(\mathrm{Tr}_{j}[U^{\dagger}OU])^{\dagger}(\mathrm{Tr}_{j}[U^{\dagger}OU])\right]\right]$
$=\sum_{j}\left[\frac{1}{d^{n}}\mathrm{Tr}[O^{\dagger}O]-\frac{1}{d^{n+1}}\mathrm{Tr}\left[(\mathrm{Tr}_{j}[O])^{\dagger}(\mathrm{Tr}_{j}[O])\right]\right].$ (16)
The second equality follows from the fact that any local unitary operation on the $j$th qubit preserves terms with $P_{j}=I$ and cannot generate such terms from those with $P_{j}\neq I$. Thus, if $U$ can be written as a product of local unitary operations as $U=\otimes_{j}U_{j}$, then $W(U^{\dagger}OU)=W(O)$ holds. In addition, since the Pauli weight is clearly preserved under swap operations, $W(O)$ is invariant under swaps as well. Therefore, the average Pauli weight is invariant under any non-entangling unitary, which can be generated as a product of single-qubit unitaries and swap operations. ∎
Proof: Additivity..
Additivity can be proved directly. By definition, the weight of Pauli strings is additive: $W(P_{1}\otimes P_{2})=W(P_{1})+W(P_{2})$. Therefore, for arbitrary normalized operators $O_{1}=\sum_{i}c_{i}P_{i}$ and $O_{2}=\sum_{j}c_{j}^{\prime}P_{j}^{\prime}$, additivity follows from the following calculation:
$W(O_{1}\otimes O_{2})$ $=\sum_{i,j}|c_{i}c_{j}^{\prime}|^{2}W(P_{i}\otimes P_{j}^{\prime})$
$=\sum_{i}|c_{i}|^{2}W(P_{i})+\sum_{j}|c_{j}^{\prime}|^{2}W(P_{j}^{\prime})$
$=W(O_{1})+W(O_{2}).$ (17)
Appendix B Quantum Fourier entropy
We provide proofs of the basic properties of the quantum Fourier entropy $S_{a}(O)$.
Proof: Faithfulness..
Faithfulness follows immediately from the definition of $S_{a}(O)$, namely, its range is from 0 to $\log_{a}(4^{n})$. In particular, when the parameter $a$ is set to 4, the maximum value of $S(O)$ is $n$. The equality $S_{a}(O)=0$ holds if and only if $O$ is a Pauli operator, since this corresponds to the case where the probability distribution is concentrated on a single element. ∎
Proof: Invariance..
Invariance follows directly from the definition of Clifford unitaries. If $U$ is a Clifford unitary and $P$ is a Pauli operator, then $U^{\dagger}PU$ is also a Pauli operator, and the mapping is injective. Therefore, when $O=\sum_{i}c_{i}P_{i}$ is conjugated by a Clifford unitary, it transforms into $U^{\dagger}OU=\sum_{i}c_{i}^{\prime}P_{i}$, where $\{c_{i}^{\prime}\}_{i}$ is a permutation of the original coefficients $\{c_{i}\}_{i}$. Since the value of the quantum Fourier entropy is invariant under permutations of the coefficients, it remains unchanged under any Clifford unitary transformation. ∎
Proof: Additivity..
Additivity can be derived through a straightforward argument. For normalized operators $O_{1}=\sum_{i}c_{i}P_{i}$ and $O_{2}=\sum_{j}c^{\prime}_{j}P^{\prime}_{j}$, their tensor product is expressed as $O_{1}\otimes O_{2}=\sum_{i,j}c_{i}c^{\prime}_{j}P_{i}\otimes P^{\prime}_{j}$. Here, $\{P_{i}\otimes P^{\prime}_{j}\}_{ij}$ forms the Pauli group on the total Hilbert space, and the coefficients $\{c_{i}c^{\prime}_{j}\}_{ij}$ satisfy the normalization condition $\sum_{ij}|c_{i}c^{\prime}_{j}|^{2}=1$. Therefore, we can obtain the following equality:
$S_{a}(O_{1}\otimes O_{2})$ $=-\sum_{i,j}|c_{i}c^{\prime}_{j}|^{2}\log|c_{i}c^{\prime}_{j}|^{2}$
$=-\sum_{i}|c_{i}|^{2}\log|c_{i}|^{2}-\sum_{j}|c^{\prime}_{j}|^{2}\log|c^{\prime}_{j}|^{2}$
$=S_{a}(O_{1})+S_{a}(O_{2}).$ (18)
Appendix C Operator complexity
In this section, we derive the operator $O_{\mathrm{Max}}$ that maximizes the operator complexity $R_{a}(O)$. We also compute the specific value of $W(O_{\mathrm{Max}})$, $S_{a}(O_{\mathrm{Max}})$, and $R_{a}(O_{\mathrm{Max}})$.
First, note that the values of these measures are independent of the complex phases of the coefficients of $O$. Thus, we define $p_{i}\coloneqq|c_{i}|^{2}$ and assume that the operator can be written as $O=\sum_{i}\sqrt{p_{i}}P_{i}$. To derive $O_{\mathrm{Max}}$, we solve the following optimization problem:
$\mathrm{maximize}\;\;R_{a}(O)=\sum_{i}p_{i}\left\{W(P_{i})-\log_{a}p_{i}\right\}.\;\;\;\left(\sum_{i}p_{i}=1,\;\;p_{i}\geq 0\right)$ (19)
We define the Lagrangian function $\mathcal{L}\coloneqq\sum_{i}p_{i}\{W(P_{i})-\log_{a}p_{i}\}-\lambda(\sum_{i}p_{i}-1)$. Then, the optimal coefficients $\{p_{i}\}$ satisfy the following conditions:
$\frac{\partial\mathcal{L}}{\partial p_{i}}=w-\log_{a}p_{i}-\log_{a}e-\lambda=0\;\;\;(w=W(P_{i})).$ (20)
From Eq. (20), we obtain $p_{i}=a^{w-\lambda}/e=:p_{w}$, where $e$ denotes the base of the natural logarithm. We note that the number of Pauli operators satisfying $W(P_{i})=w$ is given by $_{n}C_{w}\cdot 3^{w}$. From the normalization condition, we have:
$\sum_{i}p_{i}=\sum_{w=0}^{n}({}_{n}C_{w}\cdot 3^{w})\frac{a^{w-\lambda}}{e}=1.$ (21)
Therefore, we obtain
$a^{-\lambda}=\frac{e}{(1+3a)^{n}},\;\;\;\;\;p_{w}=\frac{a^{w}}{(1+3a)^{n}}.$ (22)
Taking into account the complex phases, the operator $O_{\mathrm{Max}}$ can be written as
$O_{\mathrm{Max}}=\sum_{w=0}^{n}\sum_{W(P_{i})=w}e^{i\theta_{i}}\sqrt{\frac{a^{w}}{(1+3a)^{n}}}P_{i}.$ (23)
Next, we provide a proof of Eq. (6) presented in the main text. The value of $W(O_{\mathrm{Max}})$ can be expressed as
$W(O_{\mathrm{Max}})=\sum_{w=0}^{n}({}_{n}C_{w}\cdot 3^{w})p_{w}w=\frac{1}{(1+3a)^{n}}\sum_{w=0}^{n}{}_{n}C_{w}\cdot(3a)^{w}w=:\frac{1}{(1+3a)^{n}}S_{n}.$ (24)
To derive Eq. (6), we prove the following equality by mathematical induction:
$S_{n}=\sum_{w=0}^{n}{}_{n}C_{w}\cdot(3a)^{w}w=3an(1+3a)^{n-1}.$ (25)
First, when $n=1$, Eq. (25) clearly holds. Next, assume that Eq. (25) holds for some natural number $n$. Under this assummption, we have
$(1+3a)S_{n}$ $=\sum_{w=0}^{n}{}_{n}C_{w}(3a)^{w}w+\sum_{w=0}^{n}{}_{n}C_{w}(3a)^{w+1}w$
$=\sum_{w=1}^{n}{}_{n}C_{w}(3a)^{w}w+\sum_{w=1}^{n+1}{}_{n}C_{w-1}(3a)^{w}(w-1)$
$=\sum_{w=1}^{n}({}_{n}C_{w}+{}_{n}C_{w-1})(3a)^{w}w+(3a)^{n+1}(n+1)-\sum_{w=1}^{n+1}{}_{n}C_{w-1}(3a)^{w}$
$=\sum_{w=1}^{n}{}_{n+1}C_{w}(3a)^{w}w+(3a)^{n+1}(n+1)-\frac{1}{n+1}\sum_{w=1}^{n+1}\frac{(n+1)!}{w!(n+1-w)!}(3a)^{w}w$
$=\sum_{w=1}^{n+1}{}_{n+1}C_{w}(3a)^{w}w-\frac{1}{n+1}\sum_{w=1}^{n+1}{}_{n+1}C_{w}(3a)^{w}w$
$=\frac{n}{n+1}S_{n+1},$ (26)
from which we obtain
$S_{n+1}=\frac{n+1}{n}(1+3a)S_{n}=3a(n+1)(1+3a)^{n}.$ (27)
Eq. (27) shows that if Eq. (25) holds for some natural number $n$, it also holds for the next integer $n+1$. Therefore, by mathematical induction, Eq. (25) holds for all positive integers $n$. Using Eq. (24), we obtain
$W(O_{\mathrm{Max}})=\frac{3a}{1+3a}n.$ (28)
The value of $S(O_{\mathrm{Max}})$ can be obtained through a similar calculation:
$S(O_{\mathrm{Max}})$ $=-\sum_{w=0}^{n}({}_{n}C_{w}\cdot 3^{w})\frac{a^{w}}{(1+3a)^{n}}\log_{a}\frac{a^{w}}{(1+3a)^{n}}$
$=-\frac{1}{(1+3a)^{n}}\sum_{w=0}^{n}{}_{n}C_{w}(3a)^{w}\left\{w-n\log_{a}(1+3a)\right\}$
$=-\frac{1}{(1+3a)^{n}}\left\{S_{n}-(1+3a)^{n}n\log_{a}(1+3a)\right\}$
$=n\log_{a}(1+3a)-\frac{3a}{1+3a}n.$ (29)
From Eqs. (28) and (C), we obtain
$R(O_{\mathrm{Max}})=W(O_{\mathrm{Max}})+S(O_{\mathrm{Max}})=n\log_{a}(1+3a).$ (30)
Appendix D Entanglement growth
In this section, we provide proofs of the fundamental properties of the entanglement growth $G_{E}(U)$. The overall structure of the proofs follows Ref. [26]. However, since the invariance property stated in Ref. [26] contains an error, we correct it and provide a revised proof.
Proof: Faithfulness..
The inequality $G_{E}(U)\geq 0$ follows from the faithfulness of the average Pauli weight. For any unitary $U$ and a weight-1 Pauli operator $O$, we have
$G_{E}(U)\geq W(U^{\dagger}OU)-W(O)=W(U^{\dagger}OU)-1\geq 0.$ (31)
Moreover, from the invariance of $W(O)$, it follows that when $U$ is a non-entangling unitary (i.e., generated from single-qubit unitaries and swap operations), we have equality $G_{E}(U)=0$. Next, we show that any unitary satisfying $G_{E}(U)=0$ can be written as a product of single-qubit unitaries and swap operations. By definition, such a unitary satisfies the relation $W(U^{\dagger}OU)=W(O)$ for any operator $O$. To proceed, we consider $O=X_{1}$, where the subscript denotes the qubit on which the Pauli operator acts. From the condition $W(U^{\dagger}X_{1}U)=W(X_{1})=1$, the following must hold:
$U^{\dagger}X_{1}U=\sum_{i}\alpha_{i}X_{i}+\sum_{i}\beta_{i}Y_{i}+\sum_{i}\gamma_{i}Z_{i},$ (32)
that is, $U^{\dagger}X_{1}U$ must be expressed solely in terms of weight-1 Pauli operators. Suppose, for contradiction, that there exists a pair of indices $(i,j)$ such that $\alpha_{i}\neq 0,\alpha_{j}\neq 0,$ and $i\neq j$. Then, squaring the right-hand side of Eq. (32) yields a term of the form $2\alpha_{i}\alpha_{j}(X_{i}\otimes X_{j})$. However, this contradicts the identity $(U^{\dagger}X_{1}U)^{2}=I$. Therefore, Eq. (32) must be rewritten as
$U^{\dagger}X_{1}U=\alpha_{i}X_{i}+\beta_{j}Y_{j}+\gamma_{k}Z_{k}.$ (33)
Additionally, if $i\neq j$, squaring the right-hand side of Eq. (33) yields a term of the form $2\alpha_{i}\beta_{j}(X_{i}\otimes Y_{j})$. Hence, Eq. (33) must be further rewritten as
$U^{\dagger}X_{1}U=\alpha_{i}X_{i}+\beta_{i}Y_{i}+\gamma_{i}Z_{i}.$ (34)
Similarly, when we take $O=Z_{1}$, we can write
$U^{\dagger}Z_{1}U=\alpha^{\prime}_{j}X_{j}+\beta^{\prime}_{j}Y_{j}+\gamma^{\prime}_{j}Z_{j}.$ (35)
If $i\neq j$, then $U^{\dagger}X_{1}U,U^{\dagger}Z_{1}U]=0$ must hold. However, this would imply $X_{1},Z_{1}]=0$, which is a contradiction. Therefore, we must have $i=j$. The same argument applies to the expansion of $U^{\dagger}Y_{1}U$. Hence, for any local operator $A_{1}$, there exists a local unitary $V_{1}$ such that
$U^{\dagger}(A_{1}\otimes I_{n-1})U$ $=(V_{1}^{\dagger}A_{i}V_{1})\otimes I_{n-1}$
$(V_{1}\mathrm{SWAP}_{1,i}U)^{\dagger}(A_{1}\otimes$ $I_{n-1})(V_{1}\mathrm{SWAP}_{1,i}U)=A_{1}\otimes I_{n-1}.$ (36)
From this relation, we can write $V_{1}\mathrm{SWAP}_{1,i}U=I_{1}\otimes V_{n-1}$, where $V_{n-1}$ is a unitary operator acting on the remaining n1n-1 qubits. By iterating this process, any unitary $U$ satisfying $G_{E}(U)=0$ can be decomposed into a product of single-qubit unitaries and swap operations, as expressed in
$U=(\mathrm{SWAP}_{1,i}V_{1})(\mathrm{SWAP}_{2,j}V_{2})\cdots.$ (37)
Therefore, the necessary and sufficient condition for $G_{E}(U)=0$ is that $U$ is a non-entangling unitary.
Proof: Invariance..
For any non-entangling unitaries $V_{1}$ and $V_{2}$, the invariance of the entanglement growth follows from the following calculation:
$G_{E}(V_{1}UV_{2})$ $=\max_{O}\left\{W(V_{2}^{\dagger}U^{\dagger}V_{1}^{\dagger}OV_{1}UV_{2})-W(O)\right\}$
$=\max_{O}\left\{W(U^{\dagger}V_{1}^{\dagger}OV_{1}U)-W(O)\right\}$
$=\max_{O}\left\{W(U^{\dagger}OU)-W(V_{1}OV_{1}^{\dagger})\right\}$
$=\max_{O}\left\{W(U^{\dagger}OU)-W(O)\right\}$
$=G_{E}(U).$ (38)
The second and fourth equalities hold due to the invariance of the average Pauli weight. ∎
Proof: Subadditivity.
We now prove the subadditivity $G_{E}(U_{1}\otimes U_{2})\leq G_{E}(U_{1})+G_{E}(U_{2})$. As a preliminary step, we first show that $G_{E}(UV)\leq G_{E}(U)+G_{E}(V)$, as follows:
$G_{E}(UV)$ $=\max_{O}\left\{W(V^{\dagger}U^{\dagger}OUV)-W(O)\right\}$
$=\max_{O}\left\{W(V^{\dagger}U^{\dagger}OUV)-W(U^{\dagger}OU)+W(U^{\dagger}OU)-W(O)\right\}$
$\leq\max_{O}\left\{W(V^{\dagger}U^{\dagger}OUV)-W(U^{\dagger}OU)\right\}+\max_{O}\left\{W(U^{\dagger}OU)-W(O)\right\}$
$=G_{E}(U)+G_{E}(V).$ (39)
Next, we prove the inequality $G_{E}(U\otimes I_{2})\leq G_{E}(U)$. Let us consider the Hilbert space $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$, where $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are $k$-qubit and (nk)(n-k)-qubit subsystems, respectively. We expand an arbitrary operator $O$ as $O=\sum_{i,j}c_{ij}P_{i}\otimes P^{\prime}_{j}$, where $\{P_{i}\}_{i}$ and $\{P_{j}^{\prime}\}_{j}$ denote the Pauli bases on $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, respectively. From the operator $O$, we define the following operators:
$O_{j}:=\frac{1}{2^{n-k}}\mathrm{Tr}_{2}[O(I_{1}\otimes O_{j})],\;\;\;B_{j}:=\frac{O_{j}}{\|O_{j}\|},$ (40)
$O_{i}:=\frac{1}{2^{k}}\mathrm{Tr}_{1}[O(P_{i}\otimes I_{2})],\;\;\;A_{i}:=\frac{O_{i}}{\|O_{i}\|}.$ (41)
Note that each $O_{i}$ acts on $\mathcal{H}_{2}$ and each $O_{j}$ acts on $\mathcal{H}_{1}$, and that $A_{i}$ and $B_{j}$ are the normalized versions of $O_{i}$ and $O_{j}$, respectively. Since $O_{j}=\sum_{i}c_{ij}P_{i}$ and $O_{i}=\sum_{j}c_{ij}P^{\prime}_{j}$, the average Pauli weight of $O$ can be expressed as
$W(O)$ $=\sum_{i,j}|c_{ij}|^{2}\left\{W(P_{i})+W(P^{\prime}_{j})\right\}$
$=\sum_{j}\|O_{j}\|^{2}W(B_{j})+\sum_{i}\|O_{i}\|^{2}W(A_{i}).$ (42)
By applying the same argument to $(U^{\dagger}\otimes I_{2})O(U\otimes I_{2})$, we obtain
$W((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))=\sum_{j}\|O_{j}\|^{2}W(U^{\dagger}B_{j}U)+\sum_{i}\|O_{i}\|^{2}W(A_{i}).$ (43)
Combining Eqs. (D) and (43), we then arrive at
$W((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))-W(O)$ $=\sum_{j}\|O_{j}\|^{2}\left\{W(U^{\dagger}B_{j}U)-W(O)\right\}$
$\leq\max_{\|B_{j}\|=1}\left\{W(U^{\dagger}B_{j}U)-W(O)\right\}$
$=G_{E}(U).$ (44)
Since this inequality holds for any operator $O$, we obtain
$G_{E}(U\otimes I_{2})\leq G_{E}(U).$ (45)
Combining Eq. (D), Eq. (45), and similarly derived inequality $G_{E}(I_{1}\otimes U)\leq G_{E}(U)$, we conclude that for any unitary $U_{1}$ acting on $\mathcal{H_{1}}$ and $U_{2}$ acting on $\mathcal{H}_{2}$, the entanglement growth is subadditive:
$G_{E}(U_{1})\leq G_{E}(U_{1})+G_{E}(U_{2}).$ (46)
Appendix E Magic growth
In this section, we prove the basic properties of the magic growth $G_{M}(U)$. The general structure of the proofs follows Ref. [26], but we correct the condition for invariance stated therein, as it contains an inaccuracy.
Proof: Faithfulness..
Faithfulness $G_{M}(U)\geq 0$ follows directly from the faithfulness of the quantum Fourier entropy. For any unitary $U$, if $O$ is a Pauli operator, then
$G_{M}(U)\geq S_{a}(U^{\dagger}OU)-S(O)=S(U^{\dagger}OU)\geq 0.$ (47)
If $U$ is a Clifford unitary, then $G_{M}(U)=0$ due to the invariance of $S_{a}(O)$. Conversely, if $U$ is a non-Clifford unitary, there exists a Pauli operator $O$ such that $U^{\dagger}OU$ is not a Pauli operator. This implies $S_{a}(U^{\dagger}OU)-S(O)>0$, and hence $G_{M}(U)>0$. Therefore, $G_{M}(U)=0$ if and only if $U$ is a Clifford unitary. ∎
Proof: Invariance..
Invariance of the magic growth can be established as follows. When both $V_{1}$ and $V_{2}$ are Clifford unitaries, we have
$G_{M}(V_{1}UV_{2})$ $=\max_{O}\left\{S(V_{2}^{\dagger}U^{\dagger}V_{1}^{\dagger}OV_{1}UV_{2})-S(O)\right\}$
$=\max_{O}\left\{S(U^{\dagger}V_{1}^{\dagger}OV_{1}U)-S(O)\right\}$
$=\max_{O}\left\{S(U^{\dagger}OU)-S(V_{1}OV_{1}^{\dagger})\right\}$
$=\max_{O}\left\{S(U^{\dagger}OU)-S(O)\right\}$
$=G_{M}(U).$ (48)
The second and fourth equalities follow from the invariance of the quantum Fourier entropy. ∎
Proof: Subadditivity..
We prove the subadditivity $G_{M}(U_{1}\otimes U_{2})\leq G_{M}(U_{1})+G_{M}(U_{2})$. The overall structure of the proof follows that of the entanglement growth case. As a first step, we show that the inequality $G_{M}(UV)\leq G_{M}(U)+G_{M}(V)$ holds as follows:
$G_{M}(UV)$ $=\max_{O}\left\{S(V^{\dagger}U^{\dagger}OUV)-S(O)\right\}$
$=\max_{O}\left\{S(V^{\dagger}U^{\dagger}OUV)-S(U^{\dagger}OU)+S(U^{\dagger}OU)-S(O)\right\}$
$\leq\max_{O}\left\{S(V^{\dagger}U^{\dagger}OUV)-S(U^{\dagger}OU)\right\}+\max_{O}\left\{S(U^{\dagger}OU)-S(O)\right\}$
$=G_{M}(U)+G_{M}(V).$ (49)
Next, we show the inequality $G_{M}(U\otimes I_{2})\leq G_{M}(U)$. Let us consider a subsystem consisting of $k$ qubits. Given an operator $O=\sum_{i,j}c_{ij}P_{i}\otimes P^{\prime}_{j}$ on the total system, we define the following operators on the subsystem:
$O_{j}:=\frac{1}{2^{n-k}}\mathrm{Tr}_{2}[O(I_{1}\otimes O_{j})],\;\;\;B_{j}:=\frac{O_{j}}{\|O_{j}\|}.$ (50)
We note that $B_{j}=\sum_{i}c_{ij}/\|O_{j}\|P_{i}$. Using this operator, we can compute $S(O)$ as
$S(O)$ $=-\sum_{i,j}|c_{ij}|^{2}\log|c_{ij}|^{2}$
$=\sum_{j}\|O_{j}\|^{2}S(B_{j})-\sum_{j}\|O_{j}\|^{2}\log\|O_{j}\|^{2}.$ (51)
By the same calculation, we obtain
$S((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))=\sum_{j}\|O_{j}\|^{2}S(U^{\dagger}B_{j}U)-\sum_{j}\|O_{j}\|^{2}\log\|O_{j}\|^{2}.$ (52)
From Eqs. (E) and (52), the following inequality holds:
$S((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))-S(O)$ $=\sum_{j}\|O_{j}\|^{2}\left\{S(U^{\dagger}B_{j}U)-S(B_{j})\right\}$
$\leq\max_{\|B_{j}\|=1}\left\{S(U^{\dagger}B_{j}U)-S(B_{j})\right\}$
$=G_{M}(U).$ (53)
Since this inequality holds for any operator $O$, we obtain
$G_{M}(U\otimes I_{2})\leq G_{M}(U).$ (54)
From Eqs. (E) and (54), together with the similarly derived inequality $G_{M}(I_{1}\otimes U)\leq G_{M}(U)$, we arrive at the subadditivity of the magic growth:
$G_{M}(U_{1}\otimes U_{2})\leq G_{M}(U_{1})+G_{M}(U_{2}).$ (55)
Appendix F Complexity growth
In this section, we prove the fundamental properties of the complexity growth. The structure of the proofs is analogous to those of the entanglement growth and the magic growth. In addition, we provide a proof of maxitivity, which is a distinctive property of $\tilde{G}(U)$.
Proof: Faithfulness..
The inequality $G(U)\geq 0$ follows directly from the faithfulness of the operator complexity. For any unitary $U$, if $O$ is a weight-1 Pauli operator, then
$G(U)\geq R(U^{\dagger}OU)-R(O)=R(U^{\dagger}OU)-1\geq 0.$ (56)
If $U$ is a non-scrambling unitary, then $G(U)=0$ due to the invariance of $R(O)$. Note that any non-scrambling unitary belongs to both the set of non-entangling unitaries and the set of Clifford unitaries. Conversely, if $U$ is not a non-entangling unitary, there exists a weight-1 Pauli operator $O$ such that $W(U^{\dagger}OU)-1+S(U^{\dagger}OU)>0$, which implies $G(U)>0$ (see the proof of the faithfulness of $W(O)$). Similarly, if $U$ is a non-Clifford unitary, there also exists a weight-1 Pauli operator $O$ such that the same inequality holds. Therefore, $G(U)=0$ if and only if $U$ is a non-scrambling unitary. By the same reasoning, the faithfulness of $\tilde{G}(U)$ can also be established.
Proof: Invariance..
Invariance can be established in the same manner as for $G_{E}(U)$ and $G_{M}(U)$. When both $V_{1}$ and $V_{2}$ are non-scrambling unitaries, we have
$G(V_{1}UV_{2})$ $=\max_{O}\left\{R(V_{2}^{\dagger}U^{\dagger}V_{1}^{\dagger}OV_{1}UV_{2})-R(O)\right\}$
$=\max_{O}\left\{R(U^{\dagger}V_{1}^{\dagger}OV_{1}U)-R(O)\right\}$
$=\max_{O}\left\{R(U^{\dagger}OU)-R(V_{1}OV_{1}^{\dagger})\right\}$
$=\max_{O}\left\{R(U^{\dagger}OU)-R(O)\right\}$
$=G(U).$ (57)
The second and fourth equalities follow from the invariance of $R(O)$. By the same calculation, we obtain $\tilde{G}(V_{1}UV_{2})=\tilde{G}(U)$.
Proof: Subadditivity..
We prove the subadditivity of $G(U)$ following the same structure as in the proofs for the entanglement growth and the magic growth. We begin by showing that the inequality $G(UV)\leq G(U)+G(V)$ holds for any unitaries $U$ and $V$:
$G(UV)$ $=\max_{O}\left\{R(V^{\dagger}U^{\dagger}OUV)-R(O)\right\}$
$=\max_{O}\left\{R(V^{\dagger}U^{\dagger}OUV)-R(U^{\dagger}OU)+R(U^{\dagger}OU)-R(O)\right\}$
$\leq\max_{O}\left\{R(V^{\dagger}U^{\dagger}OUV)-R(U^{\dagger}OU)\right\}+\max_{O}\left\{R(U^{\dagger}OU)-R(O)\right\}$
$=G(U)+G(V).$ (58)
Next, we prove the inequality $G(U\otimes I_{2})\leq G(U)$. We expand an arbitrary operator $O$ as $O=\sum_{i,j}c_{ij}P_{i}\otimes P^{\prime}_{j}$. Following the definitions used in the proofs of the entanglement growth and the magic growth, we introduce the following operators:
$O_{j}:=\frac{1}{2^{n-k}}\mathrm{Tr}_{2}[O(I_{1}\otimes O_{j})],\;\;\;B_{j}:=\frac{O_{j}}{\|O_{j}\|},$ (59)
$O_{i}:=\frac{1}{2^{k}}\mathrm{Tr}_{1}[O(P_{i}\otimes I_{2})],\;\;\;A_{i}:=\frac{O_{i}}{\|O_{i}\|}.$ (60)
By a similar calculation to those in the previous proofs, we obtain the following expression:
$R(O)$ $=W(O)+S(O)$
$=\sum_{j}\|O_{j}\|^{2}W(B_{j})+\sum_{i}\|O_{i}\|^{2}W(A_{i})+\sum_{j}\|O_{j}\|^{2}S(B_{j})-\sum_{j}\|O_{j}\|^{2}\log\|O_{j}\|^{2}.$ (61)
Applying this calculation to the operator $((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))$, we obtain
$R((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))$
$=\sum_{j}\|O_{j}\|^{2}W(U^{\dagger}B_{j}U)+\sum_{i}\|O_{i}\|^{2}W(A_{i})+\sum_{j}\|O_{j}\|^{2}S(U^{\dagger}B_{j}U)-\sum_{j}\|O_{j}\|^{2}\log\|O_{j}\|^{2}.$ (62)
From Eqs. (F) and (F), we obtain
$R((U^{\dagger}\otimes I_{2})O(U\otimes I_{2}))-R(O)$ $=\sum_{j}\|O_{j}\|^{2}\left\{W(U^{\dagger}OU)-W(O)+S(U^{\dagger}B_{j}U)-S(B_{j})\right\}$
$\leq\max_{\|B_{j}\|=1}\left\{W(U^{\dagger}OU)+S(U^{\dagger}OU)-W(O)-S(O)\right\}$
$=G(U).$ (63)
Since this inequality holds for any operator $O$, we obtain
$G(U\otimes I_{2})\leq G(U).$ (64)
Combining Eqs. (F) and (64) with the similarly derived inequality $G(I_{1}\otimes U)\leq G(U)$, we complete the proof of the subadditivity of $G(U)$:
$G(U_{1}\otimes U_{2})\leq G(U_{1})+G(U_{2}).$ (65)
Proof: Maxitivity..
Maxitivity of $\tilde{G}(U)$ can be established through the following calculation:
$\tilde{G}(U_{1}\otimes U_{2})$ $=\max_{W(O)=1,S(O)=0}\left\{R((U_{1}^{\dagger}\otimes U_{2}^{\dagger})O(U_{1}\otimes U_{2}))-1\right\}$
$=\max\left\{\max_{W(O_{1})=1,S(O_{1})=0}\{R(U_{1}^{\dagger}OU_{1})-1\},\max_{W(O_{2})=1,S(O_{2})=0}\{R(U_{2}^{\dagger}OU_{2})-1\}\right\}$
$=\max\left\{\tilde{G}(U_{1}),\tilde{G}(U_{2})\right\}.$ (66)
Here, we note that an operator $O$ satisfies $W(O)=1$ and $S(O)=0$ if and only if $O$ is a weight-1 Pauli operator. The second equality follows from the faithfulness and additivity of the operator complexity.