### v2

# Unified quantification of entanglement and magic in information scrambling and their trade-off relation

Abstract
Entanglement and magic are among the most fundamental properties unique to quantum systems. Each quantity captures a different aspect of non-classical behavior, and each can be regarded as a resource within its own operational setting. However, the interrelation between them has not yet been fully clarified, and whether a more fundamental measure exists remains an open question. Addressing these issues is essential for deepening our understanding of quantumness. In this study, we establish a unified resource theory of information scrambling, consisting of two types: entanglement scrambling and magic scrambling. We introduce a measure that jointly characterizes both types of scrambling. This unified approach reveals a rigorous trade-off relation between entanglement and magic scrambling, as the exact maximum value of the proposed measure can be derived analytically. Furthermore, we quantify the scrambling capability of unitary transformations in terms of their ability to amplify this measure. Our work provides insights into the connection between entanglement and magic that extend beyond the context of information scrambling.
Introduction.—Quantum systems exhibit a variety of non-classical and counterintuitive behaviors. Understanding the origins of such quantumness remains one of the most fundamental challenges in quantum mechanics. Quantum resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] provide a powerful framework to capture and formalize the essential features of quantum systems (see Ref. [26] for a review). In a typical framework, a fixed set of operations, known as free operations, is specified in advance, and any property that cannot be generated using only these operations is regarded as a resource. Constructing a resource theory for a given quantum property enables both its characterization and quantitative analysis.
Entanglement and magic are key characteristics unique to quantum systems. Each of these properties embodies quantumness, but they characterize different aspects, and their relationship remains unclear in several respects. For instance, the GHZ state is highly entangled and exhibits strong quantum behavior. However, it provides no quantum advantage from the perspective of stabilizer computation, as it can be generated solely by Clifford unitaries. Conversely, the concatenation of T states exhibits large magic and enables quantum computational supremacy, even though it contains no entanglement. From these observations, one natural question arises: is there any other crucial quantity that represents quantumness? Recent studies have explored the overlap between entanglement and magic [27, 28, 29, 30, 31]. Importantly, neither resource on its own can fully account for the source of quantum advantage in certain contexts, such as stabilizer tensor networks [32]. The existence of a unified resource measure that incorporates both resources has been suggested, although its explicit form has not yet been specified [32, 33].
In this study, we establish a unified resource-theoretic framework of information scrambling and derive the rigorous trade-off relation between entanglement and magic within this framework. Information scrambling refers to the phenomenon in which information initially localized in a small subsystem becomes distributed throughout the entire system [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. In the resource-theoretic context, scrambling is categorized into two distinct types based on their underlying mechanisms [47, 48]: entanglement scrambling and magic scrambling (see FIG. 1). Each type can be regarded as a resource under its own set of free operations. However, treating them separately fails to reveal how they may influence or complement one another. We define measures that treat both entanglement scrambling and magic scrambling in parallel. Furthermore, we emphasize that entanglement scrambling and magic scrambling cannot be simultaneously maximized. To clarify their interrelationship, we introduce a measure, termed operator complexity, which captures both types of scrambling in a unified manner. As our most important result, we analytically derive the exact maximum value of the operator complexity, thereby revealing a fundamental trade-off inherent to information scrambling. This discovery is expected to advance our understanding of how entanglement and magic are connected beyond the scope of information scrambling. In addition, we propose a unified scrambling measure as an indicator of the ability to amplify the operator complexity. These quantities satisfy several essential properties of resource monotones, ensuring the consistency and soundness of our framework. Proofs for most propositions are provided in [49] and the End Matter.
Figure 1: Schematic diagrams illustrating two types of information scrambling. (a) Entanglement scrambling corresponds to transformations that increase the number of non-identity operators in Pauli strings. (b) Magic scrambling corresponds to transformations that increase the number of terms required in the Pauli basis expansion.
Scrambling of operators.—Information scrambling can be defined as the ability of unitary transformations to amplify the complexity of operators. Entanglement scrambling and magic scrambling, which are different types of scrambling arising from distinct mechanisms, can be quantified using different measures of the complexity of operators. In this section, we introduce two measures of operators that correspond to each scrambling in the context of resource theory. Furthermore, we define a unified measure that takes into account both scrambling and derive its rigorous maximum value. This clarifies the fundamental trade-off between the two types of scrambling.
We consider an $n$-qubit system in the $2^{n}$-dimensional Hilbert space $\mathcal{H}$. Let $O$ be a normalized operator acting on $\mathcal{H}$, which can be expanded in the Pauli basis as $O=\sum_{i}c_{i}P_{i}$. Here, $\{P_{i}\}_{i}$ denotes the subset of the Pauli group on $\mathcal{H}$ consisting of elements with coefficients equal to +1+1. The coefficients satisfy the condition $\sum_{i}|c_{i}|^{2}=1$, ensuring that the norm of $O$ is normalized to unity.
Firstly, we focus on entanglement scrambling. As illustrated in FIG. 1, entanglement scrambling corresponds to transformations that increase the number of non-identity operators in the Pauli strings. To quantify this spreading, we introduce the average Pauli weight [47] of the normalized operator $O$, defined as
$W(O)\coloneqq\sum_{i}|c_{i}|^{2}W(P_{i}).$ (1)
For any Pauli operator $P$, $W(P)$ denotes the number of non-identity operators composing the Pauli string. For any normalized operator $O$, the average Pauli weight has the following basic properties:
  1. 1.
    Faithfulness: $W(O)\geq 0$, and $W(O)=0$ if and only if $O$ is the identity operator.
  2. $\tilde{1}$.
    Faithfulness (traceless case): $W(O)\geq 1$ for any traceless $O$, and $W(O)=1$ if and only if $O$ can be expanded using only weight-1 Pauli operators.
  3. 2.
    Invariance: $W(U^{\dagger}OU)=W(O)$ for any non-entangling unitary $U$.
  4. 3.
    Additivity: $W(O_{1}\otimes O_{2})=W(O_{1})+W(O_{2})$ for any normalized operators $O_{1}$ and $O_{2}$.
Here, a non-entangling unitary refers to a unitary that can be generated from single-qubit unitaries and swap operations.
Next, we consider the quantification of magic scrambling, which corresponds to an increase in the number of terms in the Pauli basis expansion of the operator (see FIG. 1). We introduce a measure of the complexity of the Pauli basis expansion, referred to as the quantum Fourier entropy [48], defined as
$S_{a}(O)\coloneqq-\sum_{i}|c_{i}|^{2}\log_{a}|c_{i}|^{2}.$ (2)
The advantages of this operator-based measure are discussed in Ref. [50]. Here, we take the base of the logarithm as a parameter $a$. If $a$ is omitted, we set $a=4$ so that the maximum value of the quantum Fourier entropy becomes $n$. This choice aligns the ranges of $W(O)$ and $S(O)$, enabling us to treat the two types of scrambling on equal footing. For any normalized operator $O$, the quantum Fourier entropy satisfies the following properties:
  1. 1.
    Faithfulness: $S_{a}(O)\geq 0$, and $S_{a}(O)=0$ if and only if $O$ is a Pauli operator.
  2. 2.
    Invariance: $S_{a}(U^{\dagger}OU)=S_{a}(O)$ for any Clifford unitary $U$.
  3. 3.
    Additivity: $S_{a}(O_{1}\otimes O_{2})=S_{a}(O_{1})+S_{a}(O_{2})$ for any normalized operators $O_{1}$ and $O_{2}$.
$W(O)$ and $S_{a}(O)$ serve as resource measures under the respective resource theories where non-entangling unitaries and Clifford unitaries are considered free operations. Here, we define a resource-theoretic measure of the complexity of operators. Since this function captures both types of scrambling, it enables a unified treatment of information scrambling:
$R_{a}(O)\coloneqq W(O)+S_{a}(O)=\sum_{i}|c_{i}|^{2}\log_{a}\frac{a^{W(P_{i})}}{|c_{i}|^{2}}.$ (3)
We refer to this measure as operator complexity. The main contribution of this study is the introduction and analysis of the operator complexity. $R_{a}(O)$ quantifies the extent to which a normalized operator simultaneously exhibits entanglement and magic scrambling. Since $W(O)$ and $S(O)$ share the same range of values, when the parameter $a$ is set to 4, Eq. (3) treats both types of scrambling equally. For any normalized operator $O$, the operator complexity satisfies the following properties:
  1. 1.
    Faithfulness: $R_{a}(O)\geq 0$, and $R_{a}(O)=0$ if and only if $O$ is the identity operator.
  2. $\tilde{1}$.
    Faithfulness (traceless case): $R_{a}(O)\geq 1$ for any traceless $O$, and $R_{a}(O)=1$ if and only if $O$ is a weight-1 Pauli operator.
  3. 2.
    Invariance: $R_{a}(U^{\dagger}OU)=R_{a}(O)$ for any non-scrambling unitary $U$.
  4. 3.
    Additivity: $R_{a}(O_{1}\otimes O_{2})=R_{a}(O_{1})+R_{a}(O_{2})$ for any normalized operators $O_{1}$ and $O_{2}$.
Here, non-scrambling unitaries are defined as the intersection of non-entangling unitaries and Clifford unitaries; that is, any such unitary can be generated from single-qubit Clifford unitaries and swap operations. These properties ensure that $R_{a}(O)$ is a valid resource monotone. They can be readily derived from the properties of the average Pauli weight and the quantum Fourier entropy.
$W(O)$ and $S_{a}(O)$ cannot be simultaneously maximized; that is, there exists a trade-off between entanglement scrambling and magic scrambling. Here, we derive the exact maximum value of the operator complexity. This result reveals a fundamental limitation inherent in information scrambling.
Since the value of the operator complexity is independent of the complex phases of the coefficients, we represent the normalized operator $O$ as $O=\sum_{i}\sqrt{p_{i}}P_{i}$. Under the constraints $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$, we solve the following optimization problem:
$\mathrm{maximize}\;\;R_{a}(O)=\sum_{i}p_{i}\left\{W(P_{i})-\log_{a}p_{i}\right\}.$ (4)
We provide the solution to this problem in the End Matter. As a result, the closed-form expression of the operator that maximizes $R_{a}(O)$ is derived as
$O_{\mathrm{Max}}=\sum_{w=0}^{n}\sum_{W(P_{i})=w}e^{i\theta_{i}}\sqrt{\frac{a^{w}}{(1+3a)^{n}}}P_{i},$ (5)
where $\{\theta_{i}\}_{i}$ are arbitrary phases. This operator exhibits optimal spreading from both scrambling perspectives. Furthermore, the values of each measure obtained from $O_{\mathrm{Max}}$ can be computed as follows:
$W(O_{\mathrm{Max}})=\frac{3a}{1+3a}n,$ (6)
$S_{a}(O_{\mathrm{Max}})=n\log_{a}(1+3a)-\frac{3a}{1+3a}n,$ (7)
$R_{a}(O_{\mathrm{Max}})=n\log_{a}(1+3a).$ (8)
See the End Matter for the details of the calculation. In particular, when the parameter $a$ is set to 4, we obtain
$R(O)\leq R(O_{\mathrm{Max}})\simeq 1.85n<2n.$ (9)
This is a key finding that stems from the operator complexity. Since the maximum values of $W(O)$ and $S(O)$ are both $n$, the operator complexity is clearly bounded above by $2n$. However, Eq. (9) reveals the upper bound strictly smaller than $2n$. This bound represents the exact maximum of the operator complexity, reflecting a fundamental constraint inherent in scrambling dynamics. Similarly, for general values of the parameter $a$, we obtain the following inequality:
$R_{a}(O)\leq R_{a}(O_{\mathrm{Max}})=n\log_{a}(1+3a)<(1+\lambda)n,$ (10)
where $\lambda=\log_{a}4$. The quantity $(1+\lambda)n$ is a trivial upper bound, since $R_{a}(O)=W(O)+\lambda S(O)$. Eq. (10) also captures the intrinsic limitation between the two types of scrambling.
Figure 2: Overview of the $W$-$S$ plane for arbitrary linear operators. A key observation is the trade-off relation between the two types of scrambling, illustrated by the solid red curve. When $W(O)$ reaches its maximum, $S_{a}(O)$ cannot, and vice versa: when $S_{a}(O)$ is maximized, $W(O)$ takes a specific, non-maximal value. The dashed curve provides a rough sketch of the general trend (possibly imprecise).
In FIG. 2, we map the several representative operators on the $W$-$S$ plane. First, any weight-$n$ Pauli operator, such as $O_{1}=Z_{1}\otimes\cdots\otimes Z_{n}$, gives the maximum value of $W(O)$, while $S_{a}(O)$ remains zero. Linear combinations of weight-$n$ Pauli operators also attain the maximum value of $W(O)$, and although they yield nonzero values of $S_{a}(O)$, they do not maximize it. For instance, the operator $O_{2}$ shown in FIG. 2 maximizes $W(O)$ and achieves $S_{a}(O_{2})=n\log_{a}3$, which is smaller than the maximum value of $S_{a}(O)$. In contrast, the uniform linear combination of all Pauli basis elements with equal amplitude, $O_{3}=1/2^{n}\sum_{i}P_{i}$, maximizes $S_{a}(O)$, while its average Pauli weight is not maximized: $W(O_{3})=3n/4$. Most importantly, $O_{\mathrm{Max}}$ is optimal with respect to the operator complexity $R_{a}(O)$, even though it does not individually maximize $W(O)$ or $S_{a}(O)$. The exact location of $O_{\mathrm{Max}}$ as a function of the parameter $a$ is also illustrated in FIG 2. In the limit $a\rightarrow 1$, $O_{\mathrm{Max}}$ converges to $O_{3}$, whereas in the limit $a\rightarrow\infty$, it converges to $O_{2}$.
Table 1: Overview of the measures for each type of information scrambling. A parallel structure emerges based on the resource-theoretic framework, with each type of scrambling associated with its own set of free operations.
  Operator   Unitary   Free unitary
  $W(O)=\sum_{i}|c_{i}|^{2}W(P_{i})$   $G_{E}(U)\coloneqq\max\left[W(U^{\dagger}OU)-W(O)\right]$   Non-entangling unitary
  $S(O)=-\sum_{i}|c_{i}|^{2}\log|c_{i}|^{2}$   $G_{M}(U)\coloneqq\max\left[S(U^{\dagger}OU)-S(O)\right]$   Clifford unitary
  $R(O)=\sum_{i}|c_{i}|^{2}(W(P_{i})-\log|c_{i}|^{2})$   $G(U)\coloneqq\max\left[R(U^{\dagger}OU)-R(O)\right]$   Non-scrambling unitary
Scrambling of unitaries.—The measures introduced in the previous section allow us to quantify information scrambling exhibited by unitary transformations. The average Pauli weight and the quantum Fourier entropy are associated with entanglement scrambling and magic scrambling, respectively. Furthermore, the operator complexity provides a unified framework for quantifying information scrambling.
We define the following measures corresponding to each type of scrambling, which we refer to as entanglement growth, magic growth, and complexity growth, respectively:
$G_{E}(U)\coloneqq\max_{\|O\|_{2}=1}\left\{W(U^{\dagger}OU)-W(O)\right\},$ (11)
$G_{M}(U)\coloneqq\max_{\|O\|_{2}=1}\left\{S(U^{\dagger}OU)-S(O)\right\},$ (12)
$G(U)\coloneqq\max_{\|O\|_{2}=1}\left\{R(U^{\dagger}OU)-R(O)\right\}.$ (13)
The definitions presented here are inspired by the idea discussed in Refs. [47, 48]. Here, we set the parameter $a$ to 4, but the definition extends naturally to general values. These functions satisfy the following fundamental properties:
  1. 1.
    Faithfulness: $G_{(E,M)}(U)\geq 0$, and $G_{(E,M)}(U)=0$ if and only if $U$ is a free unitary.
  2. 2.
    Invariance: $G_{(E,M)}(V_{1}UV_{2})=G_{(E,M)}(U)$ for any free unitaries $V_{1},V_{2}$.
  3. 3.
    Subadditivity: $G_{(E,M)}(U_{1}\otimes U_{2})\leq G_{(E,M)}(U_{1})+G_{(E,M)}(U_{2})$ for any unitaries $U_{1}$ and $U_{2}$.
Non-entangling unitaries, Clifford unitaries, and non-scrambling unitaries constitute the set of free unitaries for $G_{E}(U)$, $G_{M}(U)$, and $G(U)$, respectively. Thus, these serve as valid measures from the perspective of resource theory.
Here, we provide an overview of the resource-theoretic framework for information scrambling (see TABLE 1.). We introduced three distinct operator-based measures; $W(O)$, $S(O)$, and $R(O)$. In particular, $R(O)$ captures both entanglement and magic scrambling. Based on the ability of unitary transformations to amplify these scrambling-related features, we defined corresponding measures of information scrambling. These measures satisfy several fundamental properties, allowing them to be treated as resources. For each type of scrambling, an appropriate set of free unitaries is identified. Taken together, the framework exhibits a parallel structure across the different scrambling types.
Finally, we introduce an additional measure of complexity growth:
$\tilde{G}(U)\coloneqq\max_{W(O)=1,S(O)=0}\left\{R(U^{\dagger}OU)-R(O)\right\}.$ (14)
This measure satisfies the following properties:
  1. 1.
    Faithfulness: $\tilde{G}(U)\geq 0$, and $\tilde{G}(U)=0$ if and only if $U$ is a non-scrambling unitary.
  2. 2.
    Invariance: $\tilde{G}(V_{1}UV_{2})=\tilde{G}(U)$ for any non-scrambling unitaries $V_{1},V_{2}$.
  3. 3.
    Maxitivity: $\tilde{G}(U_{1}\otimes U_{2})=\max\left\{\tilde{G}(U_{1}),\tilde{G}(U_{2})\right\}$ for any unitaries $U_{1}$ and $U_{2}$.
These support the validity of $\tilde{G}(U)$ as a resource monotone. In particular, maxitivity, which is a stronger property than subadditivity, is a distinctive feature of $\tilde{G}(U)$. $\tilde{G}(U)$ can be regarded as a more suitable measure for quantifying the information scrambling of a localized unitary. For instance, consider a localized unitary of the form $U=U_{1}\otimes I_{2}$. In this case, $\tilde{G}(U)=\tilde{G}(U_{1})$ can accurately evaluate the scrambling capability, whereas $G(U)\leq G(U_{1})$ may dilute it. In addition, since the set of operators satisfying $W(O)=1$ and $S(O)=0$ coincides with the set of weight-1 Pauli operators, $\tilde{G}(U)$ has the advantage of being numerically computable. Taken together, these observations suggest that $\tilde{G}(U)$ is a promising candidate for a unified measure of information scrambling.
Conclusion.—In this study, we constructed a unified framework of resource theory for information scrambling. In particular, we introduced the concept of operator complexity, which accounts both entanglement scrambling and magic scrambling. As a main result, we derived the exact maximum value of this measure, revealing a trade-off relation inherent in information scrambling. In addition, we quantified the scrambling capability intrinsic to unitary transformations as the amplification of such measures. For the complexity growth, we also introduced an alternative definition, which serves as another plausible resource measure. These measures exhibit a parallel structure and satisfy the properties required of resource monotones.
As for future directions, developing efficient methods to compute concrete values of scrambling measures will be an important task. Furthermore, our framework can likely be extended to non-unitary channels, allowing the quantification of their scrambling capability. We expect that this work will deepen the understanding of the fundamental nature of information scrambling. Moreover, our analysis offers perspectives on the connection between entanglement and magic that go beyond the scope of scrambling itself. We hope that this study will inspire further developments across a broad range of topics in quantum physics.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP24KJ0960, JP23K24915 and JP24K03008.
Appendix A Operator complexity
In this section, we derive the operator $O_{\mathrm{Max}}$ that maximizes the operator complexity $R_{a}(O)$. We also compute the specific value of $W(O_{\mathrm{Max}})$, $S_{a}(O_{\mathrm{Max}})$, and $R_{a}(O_{\mathrm{Max}})$.
First, note that the values of these measures are independent of the complex phases of the coefficients of $O$. Thus, we define $p_{i}\coloneqq|c_{i}|^{2}$ and assume that the operator can be written as $O=\sum_{i}\sqrt{p_{i}}P_{i}$. To derive $O_{\mathrm{Max}}$, we solve the following optimization problem:
$\mathrm{maximize}\;\;R_{a}(O)=\sum_{i}p_{i}\left\{W(P_{i})-\log_{a}p_{i}\right\}.\;\;\;\left(\sum_{i}p_{i}=1,\;\;p_{i}\geq 0\right)$ (15)
We define the Lagrangian function $\mathcal{L}\coloneqq\sum_{i}p_{i}\{W(P_{i})-\log_{a}p_{i}\}-\lambda(\sum_{i}p_{i}-1)$. Then, the optimal coefficients $\{p_{i}\}$ satisfy the following conditions:
$\frac{\partial\mathcal{L}}{\partial p_{i}}=w-\log_{a}p_{i}-\log_{a}e-\lambda=0\;\;\;(w=W(P_{i})).$ (16)
From Eq. (16), we obtain $p_{i}=a^{w-\lambda}/e=:p_{w}$, where $e$ denotes the base of the natural logarithm. We note that the number of Pauli operators satisfying $W(P_{i})=w$ is given by $\tbinom{n}{w}\cdot 3^{w}$. From the normalization condition, we have:
$\sum_{i}p_{i}=\sum_{w=0}^{n}\binom{n}{w}3^{w}\frac{a^{w-\lambda}}{e}=1.$ (17)
Therefore, we obtain
$a^{-\lambda}=\frac{e}{(1+3a)^{n}},\;\;\;\;\;p_{w}=\frac{a^{w}}{(1+3a)^{n}}.$ (18)
Taking into account the complex phases, the operator $O_{\mathrm{Max}}$ can be written as
$O_{\mathrm{Max}}=\sum_{w=0}^{n}\sum_{W(P_{i})=w}e^{i\theta_{i}}\sqrt{\frac{a^{w}}{(1+3a)^{n}}}P_{i}.$ (19)
Next, we provide a proof of Eq. (6) presented in the main text. The value of $W(O_{\mathrm{Max}})$ can be expressed as
$W(O_{\mathrm{Max}})=\sum_{w=0}^{n}\binom{n}{w}3^{w}p_{w}w=\frac{1}{(1+3a)^{n}}\sum_{w=0}^{n}\binom{n}{w}(3a)^{w}w=:\frac{1}{(1+3a)^{n}}S_{n}.$ (20)
To derive Eq. (6), we prove the following equality by mathematical induction:
$S_{n}=\sum_{w=0}^{n}\binom{n}{w}(3a)^{w}w=3an(1+3a)^{n-1}.$ (21)
First, when $n=1$, Eq. (21) clearly holds. Next, assume that Eq. (21) holds for some natural number $n$. Under this assummption, we have
$(1+3a)S_{n}$ $=\sum_{w=0}^{n}\binom{n}{w}(3a)^{w}w+\sum_{w=0}^{n}\binom{n}{w}(3a)^{w+1}w$
$=\sum_{w=1}^{n}\binom{n}{w}(3a)^{w}w+\sum_{w=1}^{n+1}\binom{n}{w-1}(3a)^{w}(w-1)$
$=\sum_{w=1}^{n}\left\{\binom{n}{w}+\binom{n}{w-1}\right\}(3a)^{w}w+(3a)^{n+1}(n+1)-\sum_{w=1}^{n+1}\binom{n}{w-1}(3a)^{w}$
$=\sum_{w=1}^{n}\binom{n+1}{w}(3a)^{w}w+(3a)^{n+1}(n+1)-\frac{1}{n+1}\sum_{w=1}^{n+1}\frac{(n+1)!}{w!(n+1-w)!}(3a)^{w}w$
$=\sum_{w=1}^{n+1}\binom{n+1}{w}(3a)^{w}w-\frac{1}{n+1}\sum_{w=1}^{n+1}\binom{n+1}{w}(3a)^{w}w$
$=\frac{n}{n+1}S_{n+1},$ (22)
from which we obtain
$S_{n+1}=\frac{n+1}{n}(1+3a)S_{n}=3a(n+1)(1+3a)^{n}.$ (23)
Eq. (23) shows that if Eq. (21) holds for some natural number $n$, it also holds for the next integer $n+1$. Therefore, by mathematical induction, Eq. (21) holds for all positive integers $n$. Using Eq. (20), we obtain
$W(O_{\mathrm{Max}})=\frac{3a}{1+3a}n.$ (24)
The value of $S(O_{\mathrm{Max}})$ can be obtained through a similar calculation:
$S(O_{\mathrm{Max}})$ $=-\sum_{w=0}^{n}\binom{n}{w}3^{w}\frac{a^{w}}{(1+3a)^{n}}\log_{a}\frac{a^{w}}{(1+3a)^{n}}$
$=-\frac{1}{(1+3a)^{n}}\sum_{w=0}^{n}\binom{n}{w}(3a)^{w}\left\{w-n\log_{a}(1+3a)\right\}$
$=-\frac{1}{(1+3a)^{n}}\left\{S_{n}-(1+3a)^{n}n\log_{a}(1+3a)\right\}$
$=n\log_{a}(1+3a)-\frac{3a}{1+3a}n.$ (25)
From Eqs. (24) and (A), we obtain
$R(O_{\mathrm{Max}})=W(O_{\mathrm{Max}})+S(O_{\mathrm{Max}})=n\log_{a}(1+3a).$ (26)
[…]